OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 4x + y = 11 x + 2y = 8 Answers (3, 2) (2, 3) (-2, -3) (-2, 3)

OpenStudy (anonymous):

x = 8 - 2y 4(8 - 2y) + y = 11 32 - 7y = 11 -7y = -21 y = 3 therefore x = 8 - 2(3) = 2 and the answer is (2,3)